# Online Continual Graph Learning

### v1

@@ VI-C Time-incremental stream
VI-D Analysis of Design Choices
@@
Impact of batch size. Regarding the dimensions of the node batches in the stream, we can compare the results of Tables II-VII with those in Tables XI-XV of Appendix E, which are obtained with larger batch sizes (from 10 to 50 and from 50 to 250). Overall, two considered sizes have somewhat similar results. However, we observe that replay-based methods, A-GEM in particular, tend to obtain lower performances with the larger batch sizes. For these methods, a smaller batch size could thus be a regularizing factor. The fact that very small batch sizes in an online setting can lead to relatively good performance is also encouraging for the future development of OCGL techniques.
@@
Impact of sampling. By performing neighborhood sampling to process node mini-batches, we lose some information that might be helpful for the classification task. We expect therefore to trade some performance to remain within the efficiency constraint. Looking at Tables XVI-XIX (Appendix E) with results obtained with full neighborhood, we see that this is indeed true for most of the cases. On the Reddit dataset the performance degradation with sampling is more pronounced, as the number of neighbors was cut more substantially. There are nonetheless cases in which neighborhood sampling does not significantly affect performance, or even proves beneficial, such as on the Amazon Computer dataset.
@@
Sensitivity to hyperparameters. In our experiments the backbone model architecture, including number of layers and hidden units, is kept fixed. We conducted an ablation study on CoraFull to assess the impact of these choices, with full results reported and discussed in Appendix G.
@@ VII Conclusions
In this paper, we introduced the Online Continual Graph Learning (OCGL) setting, providing a formal framework that bridges the gap between the Continual Graph Learning and Online Continual Learning literature. This formulation establishes a foundation for studying graph-based learning in streaming environments where data arrives incrementally at the node level. Building on this, we raised the key issue of neighborhood expansion, proposing neighborhood sampling as a straightforward solution to bound the computational cost of training on each mini-batch even as the graph grows through time. We adapted six node classification datasets to align with the proposed OCGL framework, constructing node streams under both class-incremental and time-incremental learning scenarios. Our evaluation of suitably adapted Continual Learning methods highlights the higher performance of replay-based methods, specifically when using tailored strategies to leverage topological information to aid experience replay, yet criticalities remain in specific graph configurations. In future works, we plan to further tackle the issue of neighborhood expansion, developing tailored strategies that can ensure computational efficiency while better addressing the catastrophic forgetting problem. We further intend to consider more diverse node stream construction and additional tasks such as link prediction.
@@ Appendix A Datasets
In the experiments for this paper, we used six node-level classification graph datasets. The CoraFull dataset [66] is a citation network where nodes represent research papers and edges indicate citations between them, with labels based on paper topics. Arxiv [67] is a larger citation network derived from arXiv papers in the Computer Science category. The Reddit dataset [53] consists of posts from different communities of the Reddit platform, where nodes represent posts, and edges connect posts commented on by the same user, forming a large interaction graph. Amazon Computer [68] is a co-purchase network, where nodes are products and edges indicate frequently co-purchased items within the computers category on Amazon. In contrast to previous dataset which are homophilous, Roman Empire [69] is an heterophilous dataset obtained from the homonymous Wikipedia page, where nodes represent words connected by syntactic relationships or adjacency in the sentence. Finally, Elliptic [70] is a Bitcoin transaction dataset, consisting of transactions connected by flows of Bitcoins. Only some of the nodes are labeled, either as licit (42,019 nodes) or illicit transaction (4,545 nodes). Summary statistics for the four graphs are reported in Table VIII.
| Dataset | CoraFull | Arxiv | Amazon Computer | Roman Empire | Elliptic | |
| # nodes | 19,793 | 169,343 | 227,853 | 13,752 | 22,662 | 203,769 |
| # edges | 130,622 | 1,166,243 | 114,615,892 | 491,722 | 32,927 | 234,355 |
| # classes | 70 | 40 | 40 | 10 | 18 | 2 |
Appendix B Neighborhood expansion
Neighborhood expansion has been identified as the main issue with the online setting for graphs. In Figure 2 of the main paper we showed the phenomenon, and we report here additional plots for all datasets and batch sizes. In particular, in Figures 6-8 we measure neighborhood expansion not only in terms of number of nodes in the -hop neighborhood, as done in Section VI of the main paper, but we additionally count the number of edges between hops, which gives us a good proxy of computational time. Interestingly, the number of edges increases more drastically than the number of nodes when incrementing the number of layers, further confirming the need for a strategy such as sampling to maintain the computational complexity bounded within the limits of the online learning setting. We do not present plots for the Roman Empire and Elliptic datasets, since their very low average degree (2.9 and 1.15 respectively) effectively eliminates the neighborhood expansion problem for a small number of layers.
Appendix C Metrics
@@
To assess the performance of the model throughout the node stream, we also perform anytime evaluation, meaning that we evaluate the model on validation nodes after training on each mini-batch [28]. This allows us to capture the performance at any point in time, and observe also graphically how the model reacts to changes in data distribution. We measure this with Average Anytime Performance (AAP) [71], which is a generalization of average incremental accuracy for the online setting. Indicating with APt the average performance after training on batch , and having batches in total, . This can be interpreted as an Area Under the Curve accuracy score [28].
Appendix D Hyperparameters
A standard grid search was performed to select training hyperparameters for the models used in all experiments. We detail here the specific search space for each of the methods used in our comparisons. Two hyperparameters are common for all techniques: the learning rate, selected in the set , and the number of passes on each batch before passing to the next one, chosen between 1 and 5. No weight decay or dropout were used. Method specific hyperparameters are reported in Table IX, and specific details can be found in the original papers. In particular, the hyperparameters of regularization methods regulate the strenght of the regularization. For LwF a new hyperparameter has been introduced to adapt to the online setting: the number of batches after which to update the teacher model. For replay based methods we consider the proportion of memories to use with respect to each training batch, as using the entire buffer like in CGL is unfeasible in an online setting. The neighbors budget for SSM (both SSM-ER and SSM-A-GEM) refers to the number of the first- and second-hop neighbors to preserve in the memory budget, which influences the effective number of example in the buffer.
@@ Appendix D Hyperparameters
| A-GEM, ER, PDGNN | memory_proportion: {1,2,3} |
@@
Appendix E Extended results
E-A Full results with larger batch size
@@
In Tables XI-XIII, we provide extended results with the same setting configurations as the one explored in Section VI of the main paper, yet with larger mini-batch size. Compared to size 10 for CoraFull, Amazon Computer and Roman Empire, and 50 for Arxiv, Reddit and Elliptic, we report here batch-sizes 50 and 250 respectively.
| Method | AP% | AAPval% | AF% |
| EWC | |||
| LwF | |||
| MAS | |||
| TWP | |||
| ER | |||
| A-GEM | |||
| PDGNN | |||
| SSM-ER | |||
| SSM-A-GEM | |||
| bare | |||
| Joint | - | - |
| Method | AP% | AAPval% | AF% |
| EWC | |||
| LwF | |||
| MAS | |||
| TWP | |||
| ER | |||
| A-GEM | |||
| PDGNN | |||
| SSM-ER | |||
| SSM-A-GEM | |||
| bare | |||
| Joint | - | - |
| Method | AP% | AAPval% | AF% |
| EWC | |||
| LwF | |||
| MAS | |||
| TWP | |||
| ER | |||
| A-GEM | |||
| PDGNN | |||
| SSM-ER | |||
| SSM-A-GEM | |||
| bare | |||
| Joint | - | - |
| Method | AP% | AAPval% | AF% |
| EWC | |||
| LwF | |||
| MAS | |||
| TWP | |||
| ER | |||
| A-GEM | |||
| PDGNN | |||
| SSM-ER | |||
| SSM-A-GEM | |||
| bare | |||
| Joint | - | - |
| Method | AP% | AAPval% | AF% |
| EWC | |||
| LwF | |||
| MAS | |||
| TWP | |||
| ER | |||
| A-GEM | |||
| PDGNN | |||
| SSM-ER | |||
| SSM-A-GEM | |||
| bare | |||
| Joint | - | - |
| Method | AP% | AAPval% | AF% |
| EWC | |||
| LwF | |||
| MAS | |||
| TWP | |||
| ER | |||
| A-GEM | |||
| PDGNN | |||
| SSM-ER | |||
| SSM-A-GEM | |||
| bare | |||
| Joint | - | - |
E-B Results with full neighborhood
We report here empirical results obtained on mini-batches containing the full neighborhoods of newly presented nodes, in contrast with the main paper where neighborhood sampling was analyzed. Although the neighboring expansion problem is present, these results provide reference performance to be compared against the more realistic setting studied in the paper. For the Reddit dataset, due to the extremely high average degree, only one layer is used here for full neighborhood. Results are reported for CoraFull, Arxiv, Reddit and Amazon Computer datasets in Tables II-V, while they are not reported for Roman Empire and Elliptic as only a handful of nodes have a higher degree than 10 used for neighborhood sampling in the main experimentss of the paper. To be consistent with the use of full neighborhood information, the results reported for SSM are obtained by using the same number of seed nodes in the buffer, but with full neighborhood information as well. On the contrary, by using heavily sparsified subgraphs for the buffer in conjunction with full neighborhood for the node stream, we observed a degradation of performance, likely due to the different normalization weighting inside the GCN layer between the stream and the memory buffer.
| Method | Batch size 10 | Batch size 50 | |||||
| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
| EWC | |||||||
| LwF | |||||||
| MAS | |||||||
| TWP | |||||||
| ER | |||||||
| A-GEM | |||||||
| SSM-ER | |||||||
| SSM-A-GEM | |||||||
| bare | |||||||
| Joint | - | - | - | - | |||
| Method | Batch size 50 | Batch size 250 | |||||
| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
| EWC | |||||||
| LwF | |||||||
| MAS | |||||||
| TWP | |||||||
| ER | |||||||
| A-GEM | |||||||
| SSM-ER | |||||||
| SSM-A-GEM | |||||||
| bare | |||||||
| Joint | - | - | - | - | |||
| Method | Batch size 50 | Batch size 250 | |||||
| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
| EWC | |||||||
| LwF | |||||||
| MAS | |||||||
| TWP | |||||||
| ER | |||||||
| A-GEM | |||||||
| SSM-ER | |||||||
| SSM-A-GEM | |||||||
| bare | |||||||
| Joint | - | - | - | - | |||
| Method | Batch size 10 | Batch size 50 | |||||
| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
| EWC | |||||||
| LwF | |||||||
| MAS | |||||||
| TWP | |||||||
| ER | |||||||
| A-GEM | |||||||
| SSM-ER | |||||||
| SSM-A-GEM | |||||||
| bare | |||||||
| Joint | - | - | - | - | |||
E-C Impact of buffer size
For completeness of our results on the impact of buffer size on model performance, we show in Figures 10-11 plots similar to Figure 3 of the main body of the paper, showing both Average Performance and Average Anytime Performance achieved by replay methods with different buffer sizes, with both possible choices of mini-batch size as well.
Appendix F Anytime evaluation plots
We show here the plots with anytime evaluation, on all datasets and with the batch size used in the main experiments for the paper. In Figures 14-18 the lines indicate AP on validation nodes after training on each mini-batch, for all considered methods. In Figures 19-24, instead, we show the performance of each method in more detail, with a breakdown of accuracy by each task.
Appendix G Ablation study on CoraFull
In our experiments, the backbone model for all Continual Learning strategies was fixed as a 2-layer GCN with 256 hidden units, following Zhang et al. [46]. Thus, we performed a small ablation study on the CoraFull dataset to assess the impact of the number of hidden units and number of GCN layers. This is conducted only with full neighborhood aggregation.
G-A Number of hidden units
Still maintaining a 2-layer GCN, we changed the number of hidden units, from the original 256 to alternatively 128 and 512. We report the results in Tables XX-XXI. We observe how overall performance are lower than with 256 hidden units, possibly as this number may have been validated in previous works. A smaller network seems to consistently damage performance metrics, while with 512 units the comparison is more mixed.
| Method | Batch size 10 | Batch size 50 | |||||
| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
| EWC | |||||||
| LwF | |||||||
| MAS | |||||||
| TWP | |||||||
| ER | |||||||
| A-GEM | |||||||
| SSM-ER | |||||||
| SSM-A-GEM | |||||||
| bare | |||||||
| Method | Batch size 10 | Batch size 50 | |||||
| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
| EWC | |||||||
| LwF | |||||||
| MAS | |||||||
| TWP | |||||||
| ER | |||||||
| A-GEM | |||||||
| SSM-ER | |||||||
| SSM-A-GEM | |||||||
| bare | |||||||
G-B Number of GCN layers
We considered here a different number of GCN layers compared to our main results, 1 and 3 specifically, as on CoraFull also the 3-hop neighborhood has a relatively limited expansion. Results are reported in Tables XXII-XXIII. We observe overall lower results specifically with 3 layers, while ER and A-GEM improve using only 1 GCN layer.
Additionally, as a sort of “0-layer” baseline, we test the usage of a 2 layer MLP, equivalent to our main backbone GCN applied using the identity as adjacency matrix. This is the same as discarding graph topology information. For this reason, SSM is not used in this case. While most results are consequently greatly reduced, ER, GEM and EWC are able to maintain some performance even without using the graph topology.
| Method | Batch size 10 | Batch size 50 | |||||
| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
| EWC | |||||||
| LwF | |||||||
| MAS | |||||||
| TWP | |||||||
| ER | |||||||
| A-GEM | |||||||
| SSM-ER | |||||||
| SSM-A-GEM | |||||||
| bare | |||||||
| Method | Batch size 10 | Batch size 50 | |||||
| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
| EWC | |||||||
| LwF | |||||||
| MAS | |||||||
| TWP | |||||||
| ER | |||||||
| A-GEM | |||||||
| SSM-ER | |||||||
| SSM-A-GEM | |||||||
| bare | |||||||
| Method | Batch size 10 | Batch size 50 | |||||
| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
| EWC | |||||||
| LwF | |||||||
| MAS | |||||||
| TWP | |||||||
| ER | |||||||
| A-GEM | |||||||
| bare | |||||||

### v2

@@ VI-D Time-incremental stream
| Method | AAP% | AP% | AF% |
| A-GEM | |||
| ER | |||
| EWC | |||
| LwF | |||
| MAS | |||
| PDGNN | |||
| SSM-A-GEM | |||
| SSM-ER | |||
| TWP | |||
| LINEAR | |||
| bare | |||
| Joint | - | - |
VI-E Heterophily
To the best of our knowledge, the issue of heterophily in CGL has been addressed only by Zhao et al. [74]. However, their setting ignores inter-task edges, effectively making the graph less heterophilous. Since that requires unrealistic supervisory information on test nodes, we consider the Roman Empire dataset in our setting, showing results in Table IX. As can be expected, the results of all considered strategies are much lower than the upper bound. This can be the effect of two factors: first, the considered graph convolution may not be the best to deal with heterophily, and second, since nodes tend to connect with nodes of different classes, many connections are inter-task, and may thus not be observed when training on a certain class. This catastrophic drift in neighborhood composition is of course due to the class-incremental setting, which itself may not be realistic for heterophilous phenomena. While Roman Empire is not a real-world dataset, since it comes from words in a Wikipedia page, our results show that training GCNs in the OCGL setting poses significant challenges due to the drift in neighborhood composition.
| Method | AAP% | AP% | AF% |
| A-GEM | |||
| ER | |||
| EWC | |||
| LwF | |||
| MAS | |||
| PDGNN | |||
| SSM-A-GEM | |||
| SSM-ER | |||
| TWP | |||
| LINEAR | |||
| bare | |||
| Joint | - | - |
VI-F LINEAR
Despite its simplicity, LINEAR reliably obtains very good results: it ranks first in four out of eight benchmarks and within the top three in all but one. The only exception is Arxiv with time-incremental stream, where LINEAR underperforms compared to other methods. In this case the distribution drift is limited, as can be inferred by the positive forgetting in Table VIII and observed in appendix Figures 14 and 21, so even the bare 2-layer GCN baseline without CL strategy performs well. Instead, LINEAR here suffers from its limited expressive power, since it only aggregates 1-hop information, and with a linear classifier. Except for this limit case, we believe that such simplicity is precisely what allows LINEAR to overperform elsewhere: it makes it more robust to forgetting, and faster to adapt. Additionally, its simplicity also results in much lower computational requirements, making it a suitable choice when compute is limited or a high update frequency is required.
VI-G Design choices and ablation results
In the previous subsections, we have discussed empirical results obtained on mini-batches with neighborhood sampling and with some fixed design choices. Ablation results with respect to some of these choices are provided in Appendix D, on the datasets CoraFull, Amazon Computer and Arxiv. We briefly discuss here the impact of these configuration decisions.
Memory buffer size. In Figure 5, we observe the results obtained by the replay methods with different buffer sizes. While overall a larger memory buffer is associated with better performance, this is significantly visible only on CoraFull, as it is a smaller dataset with a large number of classes, and therefore a larger number of samples is required to be representative enough. The need for an increased buffer size is apparent for SSM-based buffers, especially for SSM-ER: as the number of examples is lower than the size of the buffer due to the storage of some neighboring nodes in the form of a sparsified subgraph, we can understand how with lower memory capacity this strategy may not be viable, potentially overfitting to a very small set of examples. PDGNN and LINEAR on the other hand show a consistently high performance even with smaller buffer size, indicating a more efficient use of memory, as it is topology aware.
@@ VI-G Design choices and ablation results
Impact of batch size. Regarding the dimensions of the node batches in the stream, we can compare the results of Tables II, V and III with those in Table XIII of Appendix D, which are obtained with larger batch sizes (from 10 to 50 and from 50 to 250). Overall, the two considered sizes have limited differences, as only on CoraFull we observe generally better performances with the larger batch size. On this dataset we also see a significant increase in the performance of regularization methods, with MAS becoming more competitive with the larger batch size. Nonetheless, the fact that very small batch sizes in an online setting can lead to relatively good performance is also encouraging for the future development of OCGL techniques.
@@
Impact of sampling. By performing neighborhood sampling to process node mini-batches, we lose some information that might be helpful for the classification task. We expect therefore to trade some performance to remain within the efficiency constraint. Looking at appendix Table XIV (Appendix D) with results obtained with full neighborhood, we see that this is indeed generally true, although to a very limited extent. There are actually cases in which neighborhood sampling does not significantly affect performance, or even proves beneficial, possibly acting as a regularizer.
@@
Sensitivity to hyperparameters. In our experiments the backbone model architecture, including number of layers and hidden units, is kept fixed. We conducted an ablation study to assess the impact of these choices, evaluating also a backbone GCN with 1 or 3 layers, or with 128 and 512 hidden units, with full results reported in Appendix D. While changing number of hidden units does not lead to particularly relevant changes, the use of 3 layers of GCN instead of 2 generally reduces performance, in agreement with literature suggesting that deeper networks are more prone to forgetting [75]. Results with 1 GCN layer are mixed, but still lower than the ones for LINEAR.
@@ VII Conclusions
In this paper, we introduced the Online Continual Graph Learning (OCGL) setting, which bridges the gap between Continual Graph Learning and Online Continual Learning. The OCGL formulation establishes a foundation for studying graph-based learning in streaming environments where data arrives sequentially, providing information at the level of nodes. A key problem that emerges in this setting is that of neighborhood expansion, which we addressed with neighborhood sampling as a straightforward solution to bound the computational and memory cost of training on each mini-batch, even as the graph grows through time. To support research in this area, we developed a benchmarking environment that adapted seven node-classification datasets to align with the proposed OCGL setting, constructing node streams under both class-incremental and time-incremental learning scenarios. Our evaluation compares nine suitably adapted methods from the CL literature along with LINEAR, a newly proposed simple and efficient baseline. The results indicate that replay-based methods perform best overall, especially when using tailored strategies to capture topological information in the memory buffer, while LINEAR achieves highly competitive results despite its simplicity, establishing it as a strong reference point for forthcoming research. In future works, we plan to further study the neighborhood expansion problem, developing tailored strategies that can ensure computational efficiency while better addressing catastrophic forgetting. We also intend to consider more diverse node stream construction and additional tasks such as link prediction.
@@ Appendix A Datasets
In the experiments for this paper, we used seven node-level classification graph datasets. The CoraFull dataset [66] is a citation network where nodes represent research papers and edges indicate citations between them, with labels based on paper topics. Arxiv [67] is a larger citation network derived from arXiv papers in the Computer Science category. The Reddit dataset [52] consists of posts from different communities of the Reddit platform, where nodes represent posts, and edges connect posts commented on by the same user, forming a large interaction graph. Amazon Computer [68] and Products [67] are co-purchase networks, where nodes are products and edges indicate frequently co-purchased items on Amazon. In contrast to previous dataset which are homophilous, Roman Empire [69] is an heterophilous dataset obtained from the homonymous Wikipedia page, where nodes represent words connected by syntactic relationships or adjacency in the sentence. Finally, Elliptic [70] is a Bitcoin transaction dataset, consisting of transactions connected by flows of Bitcoins. Only some of the nodes are labeled, either as licit (42,019 nodes) or illicit transaction (4,545 nodes). Summary statistics for the four graphs are reported in Table X.
| Dataset | CoraFull | Arxiv | Amazon Computer | Products | Roman Empire | Elliptic | |
| # nodes | 19,793 | 169,343 | 227,853 | 13,752 | 2,449,028 | 22,662 | 203,769 |
| # edges | 130,622 | 1,166,243 | 114,615,892 | 491,722 | 61,859,036 | 32,927 | 234,355 |
| # classes | 70 | 40 | 40 | 10 | 46 | 18 | 2 |
Appendix B Metrics
@@
More importantly, to assess the performance of the model throughout the node stream, we also perform anytime evaluation, meaning that we evaluate the model on validation nodes after training on each mini-batch [25]. This allows us to capture the performance at any point in time, and observe also graphically how the model reacts to changes in data distribution. We measure this with Average Anytime Performance (AAP) [71], which is a generalization of average incremental accuracy for the online setting. Indicating with APt the average performance after training on batch , and having batches in total, . This can be interpreted as an Area Under the Curve accuracy score [25].
Appendix C Hyperparameters
A standard grid search was performed to select training hyperparameters for the models used in all experiments. We detail here the specific search space for each of the methods used in our comparisons. Two hyperparameters are common for all techniques: the learning rate, selected in the set , and the number of passes on each batch before passing to the next one, chosen between 1 and 5. No weight decay or dropout were used. Method specific hyperparameters are reported in Table XI, and specific details can be found in the original papers. In particular, the hyperparameters of regularization methods regulate the strength of the regularization. For LwF a new hyperparameter has been introduced to adapt to the online setting: the number of batches after which to update the teacher model. For replay based methods we consider the proportion of memories to use with respect to each training batch, as using the entire buffer like in CGL is unfeasible in an online setting. The neighbors budget for SSM (both SSM-ER and SSM-A-GEM) refers to the number of the first- and second-hop neighbors to preserve in the memory budget, which influences the effective number of example in the buffer.
@@ Appendix C Hyperparameters
| A-GEM, ER, PDGNN, LINEAR | memory_proportion: {1,2,3} |
@@
Appendix D Ablation Study
In the main experiments of the papers, some hyperparameter and design choices, such as batch sizes and the structure of the GCN model, are kept fixed. In this section we provide results to assess the impact of these choices, for three of the considered benchmarks: CoraFull, Amazon Computer and Arxiv (in class-incremental stream setting). For ease of reference, we report here their aggregated results from the tables in the main paper into Table XII.
| CoraFull | Amazon Computer | Arxiv | |||||||
| Method | AAP% | AP% | AF% | AAP% | AP% | AF% | AAP% | AP% | AF% |
| A-GEM | |||||||||
| ER | |||||||||
| EWC | |||||||||
| LwF | |||||||||
| MAS | |||||||||
| PDGNN | |||||||||
| SSM-A-GEM | |||||||||
| SSM-ER | |||||||||
| TWP | |||||||||
| LINEAR | |||||||||
| bare | |||||||||
D-A Full results with larger batch size
@@
In Table XIII, we provide extended results with the same setting configurations as the one explored in Section VI of the main paper, yet with larger mini-batch size. Compared to size 10 for CoraFull and Amazon Computer, and 50 for Arxiv, here we use mini-batches of 50 and 250 nodes respectively. While on CoraFull and Arxiv the differences compared to the main setting are limited, on CoraFull we observe generally better performances (except for ER), with significantly higher results for regularization methods EWC and MAS.
| CoraFull | Amazon Computer | Arxiv | |||||||
| Method | AAP% | AP% | AF% | AAP% | AP% | AF% | AAP% | AP% | AF% |
| A-GEM | |||||||||
| ER | |||||||||
| EWC | |||||||||
| LwF | |||||||||
| MAS | |||||||||
| PDGNN | |||||||||
| SSM-A-GEM | |||||||||
| SSM-ER | |||||||||
| TWP | |||||||||
| LINEAR | |||||||||
| bare | |||||||||
D-B Results with full neighborhood
As one of the main requirements for the OCGL setting is to keep a constant computational footprint, in the experiments for the paper we adopt neighborhood sampling to tame the neighborhood expansion issue. Yet, we expect that this may come as a tradeoff with performance, since some information is thus discarded. To assess this, we report in Table XIV results obtained with mini-batches containing the full neighborhoods of newly presented nodes. We observe nonetheless how these results are in some cases lower than those with neighborhood sampling, suggesting that it could potentially act as a regularizer, and confirming the goodness of this efficient choice for the OCGL setting.
| CoraFull | Amazon Computer | Arxiv | |||||||
| Method | AAP% | AP% | AF% | AAP% | AP% | AF% | AAP% | AP% | AF% |
| A-GEM | |||||||||
| ER | |||||||||
| EWC | |||||||||
| LwF | |||||||||
| MAS | |||||||||
| PDGNN | |||||||||
| SSM-A-GEM | |||||||||
| SSM-ER | |||||||||
| TWP | |||||||||
| LINEAR | |||||||||
| bare | |||||||||
D-C Number of hidden units
Still maintaining a 2-layer GCN as in the main experiments, we changed here the number of hidden units, from the original 256 to alternatively 512 and 128. We report the results in Tables XV and XVI. LINEAR is not reported as it is a linear model.
| CoraFull | Amazon Computer | Arxiv | |||||||
| Method | AAP% | AP% | AF% | AAP% | AP% | AF% | AAP% | AP% | AF% |
| A-GEM | |||||||||
| ER | |||||||||
| EWC | |||||||||
| LwF | |||||||||
| MAS | |||||||||
| PDGNN | |||||||||
| SSM-A-GEM | |||||||||
| SSM-ER | |||||||||
| TWP | |||||||||
| bare | |||||||||
| CoraFull | Amazon Computer | Arxiv | |||||||
| Method | AAP% | AP% | AF% | AAP% | AP% | AF% | AAP% | AP% | AF% |
| A-GEM | |||||||||
| ER | |||||||||
| EWC | |||||||||
| LwF | |||||||||
| MAS | |||||||||
| PDGNN | |||||||||
| SSM-A-GEM | |||||||||
| SSM-ER | |||||||||
| TWP | |||||||||
| bare | |||||||||
D-D Number of GCN layers
We considered here a different number of GCN layers compared to the 2 our main results, 3 and 1 specifically, keeping the number of hidden units fixed to 256. Results are reported in Tables XVII and XVIII. LINEAR is not reported as it is a linear model. We observe overall lower results with 3 layers, in accordance with literature suggesting that deeper networks are more prone to forgetting [75]. With only 1 layers results are mixed, with also some cases of improved performance over the 2-layer network.
| CoraFull | Amazon Computer | Arxiv | |||||||
| Method | AAP% | AP% | AF% | AAP% | AP% | AF% | AAP% | AP% | AF% |
| A-GEM | |||||||||
| ER | |||||||||
| EWC | |||||||||
| LwF | |||||||||
| MAS | |||||||||
| PDGNN | |||||||||
| SSM-A-GEM | |||||||||
| SSM-ER | |||||||||
| TWP | |||||||||
| bare | |||||||||
| CoraFull | Amazon Computer | Arxiv | |||||||
| Method | AAP% | AP% | AF% | AAP% | AP% | AF% | AAP% | AP% | AF% |
| A-GEM | |||||||||
| ER | |||||||||
| EWC | |||||||||
| LwF | |||||||||
| MAS | |||||||||
| PDGNN | |||||||||
| SSM-A-GEM | |||||||||
| SSM-ER | |||||||||
| TWP | |||||||||
| bare | |||||||||
Appendix E Performance visualization
E-A Impact of buffer size
For completeness of our results on the impact of buffer size on model performance for replay methods, for which Figure 5 of the main body of the paper shows AAP, we show here in Figure 6 the impact on AP.
E-B Anytime evaluation plots
We show here the plots with anytime evaluation, on all datasets and with the settings used in the main experiments for the paper. In Figures 8-14 the lines indicate AP on validation nodes after training on each mini-batch, for all considered methods. We highlight the boundaries between tasks, the threshold up to which hyperparameter selection is performed, and the upper bound of joint training up to the current task. We note how it is natural and expected that accuracy tends to decrease with the batch index, as new classes are introduced and the classification task gets increasingly complex. In general, we observe that while the performance of regularization methods tends to decrease smoothly, the performance of replay-based methods shows much higher variations. In Figures 15-22, instead, we show the performance of each method in more detail, with a breakdown of performance by each task.